OpenStudy (anonymous):

Solve log3x+log9=0. round to nearest hundredth if necessary a..33 b..04 c.3 d.27

OpenStudy (freckles):

\[\log(3x)+\log(9)=0 ?\]

OpenStudy (anonymous):

how do i solve

OpenStudy (freckles):

If that is the right expression use that property I mentioned earlier log(a)+log(b)=log(ab) ----------- log(3x)+log(9)=log(3x*9)=log(27x) So you need to solve the following for x: log(27x)=0

OpenStudy (anonymous):

so multiply log(3*9)

OpenStudy (freckles):

Yes but you missed the x inside there log(27x)=0 when x= what?

OpenStudy (freckles):

I assume the base is 10 so we have \[\log_{10}(27x)=0\] Do you know how to solve for x? do 10^ on both sides

OpenStudy (anonymous):

3

OpenStudy (freckles):

\[10^{\log _{10}(27x)}=10^0\] \[27x=1\] Can you finish this ?

OpenStudy (anonymous):

yea

OpenStudy (freckles):

cool! :)